Date
Name
Other Observers
Instrument
Hours Observed
Hours Lost
Reason
Observing Program
09/23/2022
Michael Fauerbach
NA
CCD
0
10
Clouds
outer bands of TS Hermine....
09/20/2022
Van Hamme
Samec
CCD
4
5
Humidity
Observing interacting eclipsing binaries. I was able to catch 0.2 d period eclipsing binary last night for nearly a full cycle.
09/18/2022
Ross Silver
Dieter Hartmann
CCD
1
9
Humidity
Calculating photometric redshift of blazars
09/17/2022
Martha Leake
none
Andor
0
9
Humidity
Asteroid photometry. Did not open because of clouds, high wind, high humidity, rain, high aerosols (which decreased). Will download darks, biases later today. Left near morning twilight.
09/15/2022
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
0
8
Wind
Asteroid photometry
09/10/2022
Todd Vaccaro
Morgan Moore, Mel Isenbarger
Andor CCD
9
0
Software/Control PC Issues
EB program. Found the telescope and dome not parked correctly and were pointing NW (Dec +58). Zenith parking kept saying vignetting, reset tele ACE, then full reset, but never changed, though the Mirror cover stopped blinking that it was trying to open and simply said open. Camera not controllable from tele computer, so ran from Cam computer, but it wasn't filling in the headers with observer or object until the end of the night when I typed new information in. AIJ missing and C drive is full (D image drive not a problem) and had trouble getting MaximDL to open images till later on. The Focus routine on tele side would do its thing, but then plot nothing for bright star field or faint, so left as is for focus, but feel it could have been much better as mnt reporting 0.6" or less most the night; I was getting 1.4 down to 1.1". Cam computer also started up with an error saying that it failed to update Windows. I couldn't get the dome cam to work, which would've been nice since at one point I thought the dome/tele weren't lined up (Az were off by 70 degrees and I saw a lot of moonlight for an hour and a dark shadow of something bottom left on images).
09/08/2022
Yong Sheng
Dieter Hartmann, Isaiah Cox
Andor CCD
1
8
Other
I have to stop observing due to my health issue after finishing the calibration frames. But I did notice a tracking issue. The images show strips instead of point sources.
09/01/2022
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
0
8
Clouds
Asteroid photometry
09/04/2022
William Keel
DragonCon Live Astronomy
Andor CCD
3
0
No problems to report
End of night outreach session. Focus 10400. Guider again very reliable.
09/02/2022
Roi Alonso
Lisa Altinier
CCD
6
1
Clouds
WD photometry, closed for about 1h due to clouds. Autoguiding and almost everything working nicely. Mirror covers stay in "closing" status forever.
09/04/2022
Clouds
End of night for outreach event - Crab, NGC 157, NGC 7469... Focus 100500. Guider is more trouble-free and routine than I've ever found it before.
08/26/2022
james Webb
Jonathon Barata
CCD
0
8
Connection Issues
My home computer failed to connect and couldn't log in. My issues, not RM issues.
08/25/2022
Michael Fauerbach
NA
CCD
9
1
Software/Control PC Issues
For the ssecond time in a row, both the camera and telescope computer were unresponsive when I logged in and the dome said 'Doors Ajar' Autofocus failed tried restart, but that did not change anything. Seeing was poor at the beginning of the night, but got better towards morning.
08/21/2022
Michael Fauerbach
NA
CCD
8
2
Clouds
Both the camera and telescope computer were unresponsive when I logged in and the dome said 'Doors Ajar' Plenty of intermittent clouds and poor seeing wtih wind at the threshold most of the night.
08/19/2022
Roi Alonso
none
ccd
0
9
Clouds
High dense clouds and dust increasing. Did not open.
08/18/2022
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
8
0
No problems to report
Asteroid photometry
08/16/2022
Ron Kaitchuck
None
CCD
3
3
Equipment Failure
Exoplanet transits. The dome rotation failed. Peter Mack fixed that but the camera only produced “bias” frames after that. My guess is the dome was not properly synced to the telescope. My targets were then low in the west, so I alerted Peter, closed the dome and went to bed.
08/15/2022
Nuria Torres-Alba
Dieter Hartman
SARA RM
9
0
No problems to report
Clemson University observations of high-z blazars. No problems to report.
08/14/2022
Tom Rutherford
None
CCD
0
8
Clouds
M31
08/13/2022
Ron Kaitchuck
none
CCD
6
3
Equipment Failure
Dome stopped turning at 320 deg AZ I was not able to recover, but clouds came in antway.
08/12/2022
Tom Rutherford
None
CCD
7
0
Software/Control PC Issues
Asteroid, M31. Autofocuser did not finish routine.
08/11/2022
Tom Rutherford
None
CCD
7
0
Equipment Failure
Asteroid, M31. Autofocuser would not operate.
08/10/2022
William Keel
3 cats
Andor CCD
9
0
No problems to report
H II regions, radio galaxies. Focus in BVR 100300. Mirror cover remains open after close command.
08/09/2022
Yong Sheng
Dieter Hartmann
Andor CCD
10
0
No problems to report
Clemson University Blazar study The observation is pretty smooth.
08/08/2022
William Keel
3 cats
Andor CCD
9
0.5
Software/Control PC Issues
H II regions, radio galaxies. Focus 102300; software lost control of focus early on, cycled power on ACE cabinet at Brian Brondel's suggestion. (This also seems to have fixed the mirror covers not reaching their open limit switches when first opening). Never did get QSI camera to finish a readout to try guiding even after the restart. Note: mirror covers did not close at end of night (no response at all to clicking the CLOSE button).
08/05/2022
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
6
2
Other
Asteroid photometry. The dust was over the limit for couple of hours.
08/04/2022
Roi Alonso
none
Andor CCD
8.5
0.1
Equipment Failure
WDs photometry. Autoguider working, all good. Dome stuck at AZ=315 (coming from higher values). Solved with dome AZ reset, and manually turning the dome in the other direction (steps of 90 degs) until AZ 300, and then linking the dome again. Mirror petals not closing.
08/02/2022
Michael Fauerbach
NA
CCD
8
0
No problems to report
Had to restart several times, MaximDL crashed not sure what happened, maybe the CCD computer needs to be rebooted? Else No issues, good night.
08/01/2022
Stephen Gillam
Jacob Romeo
Andor Camera
8
2
Clouds
rOAp stars. One technical problem occurred. The dome rotation failed when the dome reached an azimuth of 313 degrees. We did an azimuth recycle and homed the dome. Note that the "home" did not work until we had moved the dome a small angle with the "GO" button.
08/01/2022
Stephen Gillam
None
Andor Camera
8
2
Clouds
ROAps. The dome stopped tracking late in the night. The dome azimuth was 314 degrees. An azimuth recycle got it going. Again, I had to move the dome manually before it would home, after the recycle. There is a bug in the RA, Dec offset GUI. When an RA offset is commanded of >60 arc seconds, or perhaps >59", the scope tracking is turned off. "IDLE" is indicated. No errors are reported. The scope can be slewed. But, it will not track when it reaches the target coordinates. This was fixed by closing down the telescope ACE node service, power-cycling the ACE control cabinet, and re-starting the ACE nodes service.
07/31/2022
Stephen Gillam
None
Andor Camera
10
0
Equipment Failure
ROAps. The dome stopped tracking late in the night. It was also not tracking when I started up at 4;15 EDT. In both cases, the dome azimuth was 314 degrees. An azimuth recycle got it going in both cases. Oddly, I had to move the dome manually before it would home, after the recycle.
07/30/2022
Stephen D. Gillam
Leila Alamos, Jacob Romeo
Andor
10
1
Equipment Failure
rOAp stars. Telescope tracked to just over 4 hours west and stopped tracking. Attempts to move the 'scope to the zenith failed with the following error. "Unable to complete slew due to equipment error."
07/29/2022
Stephen Gillam
Leila Alamos, Laina Talman-Bogusta, Jacob Romeo
Andor Camera
10
0
Software/Control PC Issues
Rapidly Oscillating Ap Stars. The Dropbox share service is not available.
07/23/2022
Michael Fauerbach
NA
CCD
6
2
Other
Dust was intermittent above the threshold value. Closed the dome a couple of times during the night.
07/22/2022
Martha Leake
none
Andor
4.5
3
Other
Dusty conditions (over 20 mcg/m3) to start, plus camera computer needed rebooting via APC rack, which was not easy to find; Brian helped. I was able early on to close the mirror covers, then waited for less dusty conditions. Opened dome 22:56 UT, closed it at 3:26 UT. Accidentally using another observer's camera control menu led to long darks that were not mine. Observed Comet and asteroid. Conditions clear, light wind, 0.6-1.5 " seeing; but dust at 23. Closed up; within 15 minutes dust reached 118.8. No problems with dome, mirror covers or pointing.
07/15/2022
Priya Gokuldass
None
CCD
0
9
Equipment Failure
Couldn't observe due to the Dome shutter failure
07/12/2022
Gary Henson
Rutherford
Andor CCD
4
0
No problems to report
Asteroid light curve Dome shutter still under repair but left open by ACE for night observing as weather clear and dry. Telescope and CCD working normally.
07/07/2022
Alex McDaniel
Dieter
Andor CCD
0
9
Equipment Failure
Dome still not functional
07/05/2022
Kurtis Williams
none
Andor
0
8
Other
Dome still not functional
06/25/2022
Ro
Dieter Hartmann
CCD
0
9
Equipment Failure
Could not observe due to the dome shelter failure.
06/23/2022
Yong Sheng
No
RM-griz
0
9
Equipment Failure
A text file on the RM computer says that RM will not be operational until Jun. 01 due to dome shelter failure.
06/07/2022
Roi Alonso
Lisa Altinier
CCD
0
8
Equipment Failure
Message in the "telescope computer" from Peter: THE SWITCHED POWER DISTRIBUTION UNIT (PDU) NEED TO BE POWER CYCLED. UNFORTUNATELY, IT WAS NOT CYCLED LAST NIGHT AND SO THE SYSTEM IS NOT USEABLE.
06/05/2022
Roi Alonso
Lisa Altinier
Andor
8
0
No problems to report
WDs photometry, with autoguiding working smoothly. First night in a long time with 0 issues (not even restarts of the ace connector client!).
06/03/2022
William Keel
2 cats
Andor CCD
9
0
No problems to report
Radio galaxies, dust in galaxies. Focus 97100, guider focus -50. QSI guide camera twice got stuck refusing to take exposures, even after one full restart.
05/31/2022
Kent Montgomery
none
CCD
0
8
Wind
Wind over 70 kph for first two hours of night so we gave up and never opened.
06/01/2022
Ken Watanabe
Nicolas Rimbert
CCD
7
0
No problems to report
Blazars
05/30/2022
Kent Montgomery
none
CCD
0
8
Humidity
Asteroid Photometry, 100% humidity, mountain was in a cloud, never opened
05/29/2022
Martha Leake
none
Andor
2.4
6.1
Humidity
Other issues: Clouds, high winds, Ace-client-gui not working, dome rotation needed homing. Waited for humidity to drop (long wait, and some telescopes never reached less than 100%). JKT was best, but neighboring telescopes' RH spikes usually indicated another wave of high humidity. Waited for high winds to calm. Followed quick restart when ace-gui fault occurred; no reoccurrence. Dome rotation activated when needed. Interesting night and an interesting object (P = 0.132 h) to track.
05/27/2022
james webb
none
CCD
0
10
Connection Issues
Blazar microvariability. My home situation prevented me from observing:(.\
05/26/2022
Alex
Dieter
Andor CCD
8.5
1.5
Software/Control PC Issues
Clemson Blazar Photo-z measurement program. during the night I needed to do a full reset and azimuth recycle. Other than that no major issues.
05/25/2022
Tomomi Otani
Julia Clark, Adrian Moraga
CCD
9.5
0
No problems to report
Hot subdwarf photometry time series. Everything was good tonight.
05/25/2022
Tomomi Otani
Julia Clark, Adrian Moraga
CCD
9.5
1
Software/Control PC Issues
Hot subdwarf photometry time series. ACE client stopped working a few times. The regular procedure worked to fix them.
05/22/2022
Roi Alonso
none
CCD
7.5
0.5
Equipment Failure
WDs photometry. QSI camera not responding to commands. Dome stopped rotating, AZ reset did not work, but a full reset did.
05/22/2022
Ron Kaitchuck
none
CCD
5
0
Software/Control PC Issues
The auto-focus routine was unable to change the focus. Fortunately the telescope was already at an acceptable focus.
05/22/2022
Todd Vaccaro
N/A
Andor CCD
7
.5
Software/Control PC Issues
Eclipsing binaries. Lost time with the crashing of ACE on tele computer just about every hour during sequences. Dome stuck beginning of night that an Az reset, home fixed.
05/21/2022
Todd Vaccaro
N/A
Andor ccd
6
1
Software/Control PC Issues
Eclipsing binaries. ACE crashed on the telescope computer almost every hour, which was rectified by simply closing it and restarting it. The telescope got stuck in the west near airmass 2.3, resetting HA and moving with the hand paddle eventually worked to free it, but lost 20 minutes. Catching the failing ACE during sequences lost me blocks of 5 minutes or so each hour. I noticed on the dome cam that the dome slit has a strip of insulation or something about a foot or two long dangling from near the top, which may fall into the scope. I have an image of it I can share if needed.
05/18/2022
Tomomi Otani
Julia Clark, Adrian Moraga
CCD
8
1
Software/Control PC Issues
Photometry time series of Hot subdwarf stars. The focus stopped working, but it started working again after we did a full restart.
05/17/2022
Tomomi Otani
Adrian Moraga, Julia Clark
Andor
8
1
Software/Control PC Issues
Photometry time series. Dome stopped rotating one time. We followed the procedures on the note on the telescope PC and the issue was fixed.
05/13/2022
Núria Torres-Albà
Dieter Hartmann
SARA RM
7
0
No problems to report
Clemson University observations of high-z blazars No problems to report
05/12/2022
Robert Berrington
none
camera
3
5
Clouds
short period variable stars.
05/10/2022
Roi Alonso
Lisa Altinier
CCD
7
1
Wind
WDs photometry, around 1h closed due to wind. No success with the autoguider, the QSI camera seems unresponsive (reading takes forever, reported temperature at 1000C and reported size 0x0 pixels).
05/09/2022
Roi Alonso
Lisa Altinier
CCD
0
8
Wind
Wind gusts above 45km/h all night, we did not open the dome.
05/07/2022
Roi Alonso
Lisa Altinier
CCD
8
0
No problems to report
WDs photometry with autoguiding. Perfect night. ACE client at telescope PC crashed twice, only needing to close and open the program.
05/06/2022
Yong
Dieter Hartmann
RM griz
10
0
Software/Control PC Issues
On the telescope PC, a window popped up multiple times, saying that ace stopped working. Not a big ideal since reopen the ace client will fix it. Restart is not even needed.
05/04/2022
B Murphy
Day, Grabarz, Pease
Andor
9
0
No problems to report
Globular Cluster photometry
05/03/2022
William Keel
3 cats
Andor CCD
9
0
No problems to report
AGN with emission-line clouds. Focus 97300, guider focus 800. Do not use the "save last image" button while another exposure is in progress - took me a half hour, two connector-client restarts and 2 dome azimuth resets to recover from that bit of genius. (I never did quite recover the setup on the CCD computer) Brian Brondel connected and tweaked the guider to remove E-W oscillations.
05/02/2022
William Keel
3 cats
Andor CCD
9.5
0
No problems to report
Finishing merger sample in Dead Quasar Survey. Unsure whether dome fully opened at start of night after talking to Brian Brondel. Doughnuts of dust and defocussed stars looked OK, as did camera view of dome slit at dawn. Focus 97400. Guider continues much more reliable with auto ROI and 2x2 binning.
04/30/2022
Roi Alonso
Lisa Altinier
CCD
8
0
No problems to report
WDs photometry with autoguider (thanks for the new tips!). Everything running smoothly, except ACE connector client at telescope PC crashing twice.
04/30/2022
Michael Fauerbach
NA
CCD
8.5
0.5
Software/Control PC Issues
Had to do two full restarts during the night plus an azimuth dome recycle.
04/29/2022
Michael Fauerbach
NA
CCD
7
0
No problems to report
Telescope GUI crashed twice, else n issues.
04/27/2022
Roi Alonso
Lisa Altinier
CCD
8
1
Clouds
WDs photometry. Tried to autoguide, with no success: the QSI focus does not move and seems unfocussed, the X-stage is locked between position 1 and 2, can move to 0 and 1, but stays in the middle when trying to go to 2. Usual problems: ACE restarts needed at telescope, and dom AZ resets.
04/27/2022
Michael Fauerbach
NA
CCD
8.5
0.5
Software/Control PC Issues
The camera computer suddenly rebooted without prior warning. When I logged back in a window poppoed up saying the the AVG anti-virus software has been updated. If something needs updating usually one gets a prompt, not this time. Evrything else went smooth.
04/26/2022
Michael Fauerbach
NA
CCD
5
5
Other
Thanks to Saida for offering up the night. I got a late start, else the night was rather uneventful. Focus was unresponsive once and when I tried to restart the node service Windows was unhappy and couldn't stop it. Thanks to ACE the problem was solved quickly.
04/21/2022
Roi Alonso
Lisa Altinier
CCD
1
9
Humidity
WDs. Brian and ACE prepared the telescope for the night (thanks!). No issues, apart from the humidity suddenly (minutes) going from 20% to 90%. Fortunately we noticed it and we closed the dome in time. It remained high the whole night.
04/21/2022
Stephen Gillam
None
ARC Camera
0
12
Equipment Failure
rOAp stars. The dome got stuck at 182 degrees. I homed it and then the dome tracked the telescope azimuth without another hitch. The telescope got "stuck" on the west limit at +3 hours HA. This was odd as I had slewed to the same hour angle just a few mins earlier. I checked the DIO96 bord and found that it was not operating. (The board was not present in the Instacal GUI.) I shut down the Telescope computer from software. Then powered off the Telescope control computer and the ACE control cabinet from the remote APC. A test of the DIO board showed "FAIL" and red lights on all its channels. I repeated the DIO board recycle procedure three times, occasionally recycling the power to the computer and ACE controls, instead of powering off and then on again. All attempts left the DIO board in the failed state.
04/19/2022
Ted von Hippel
none
none
0
0
No problems to report
Offered the telescope to SARA-Board yesterday but no takers.
04/13/2022
Stephen Gillam
None
Andor Camera
10
10
Wind
ROAp stars.
04/10/2022
Roi Alonso
Lisa Altinier
CCD
8.5
0.5
Equipment Failure
WDs photometry. Clouds passing all night. Typical problems with typical solutions (dome not rotating, doors ajar, ace connector resets).
04/09/2022
Yong Sheng
Dieter Hartmann
RM grizR HAlpha
11
0
Equipment Failure
Clemson Blazar study. The telescope reached the limit and couldn't slew. But it was fixed later.
04/08/2022
Martha Leake
none
Andor
9
0
No problems to report
All worked well. I did a quick restart near midnight UT and a dome azimuth restart; otherwise the dome worked and observing went smoothly. Not much dust; in the "green" all night.
04/08/2022
Brian
Students
Andor
8
1
Software/Control PC Issues
Globular cluster photometry. Problem encountered: Main Focus was not responding ... Solution: Reboot ACE cabinet. Problem encountered: Dome not moving ... Solution: Do an Azimuth reset per instructions on the desktop. (And waiting 2 minutes.) Dome opened and closed with no issue.
04/06/2022
William Keel
not even time for cats to arrive
Andor CCD
0
9
Equipment Failure
Tried opening dome per Brian Murphy's note ,and it stuck slightly open. With Peter and Brian at ACE, decided to wrangle it closed and leave it there for safety.
04/04/2022
Todd Hillwig
None.
Andor CCD
0
10
Clouds
None.
04/03/2022
Michael Fauerbach
NA
CCD
0
9
Clouds
Never opened. Clouds, high humidity, temperature around freezing.
04/02/2022
Michael Fauerbach
NA
CCD
0.1
9
Clouds
Clouds and high humidity early on. Around 0100UT sky was clear and humidity had dropped to acceptable levels. I noticed other observatories had opened, so I opened up, too. Before I was able to focus the scope, humidity rose again and I had to close. Was open for about 5 minutes.
03/26/2022
Martha A. Leake
none
Andor
0
12
Equipment Failure
Dome problems shutdown. Also today, high humidity and rain. Will not open.
03/24/2022
Alex
Dieter
Andor CCD
0
12
Equipment Failure
JKT Shutdown due to dome issue
03/22/2022
Roi Alonso
Lisa Altinier
CCD
0
9
Humidity
High humidity and temperatures bellow 0 C, with risk of ice. We did not open.
03/15/2022
Ted von Hippel
none
imaging
0
12
Other
The dome was iced over so no one would have been able to observe anyway. That having been said, because of the dome shutter problem I tried to give the night away.
03/14/2022
Roi Alonso
none
CCD
0
9
Snow
Bad weather, ICE (whenever the weather improves, please confirm it is safe to open the dome!, e.g., checking if other telescopes at the observatory are opening their domes)
03/11/2022
Yong Sheng
Yong Sheng, Dieter Hartmann
RM griz
8
2
Software/Control PC Issues
When beginning: dome stuck at "Dome is opening". Fixed it by doing full restart and cycling the Shutter Door Then the doom started vignetting. Tried all my troubleshooting methods, nothing worked, so i have to stop.
03/10/2022
Todd Vaccaro
Robert Berrington
Andor
3
6.5
Software/Control PC Issues
Eclipsing binary program. Unable to get eve flats due to dome issues (AZ resets, ACE/CCD connection required pc reboots, and shutter was having problems opening all the way). Half way through the night, and unsure if shutter open all the way, I couldn't close following instructions on the Notepad, but Brian Murphy was able to after hitting the big red X button on Dome Control, which isn't mentioned in Notepad, so we closed for the rest of the night as I was able to get the eclipse I was looking for earlier than expected.
03/07/2022
Kurtis Williams
none
Andor
0
0
Other
Addendum: Ace Connector Client GUI on telescope computer crashed twice during the night (Windows "gui stopped working" error). Did not affect observing - telescope stayed tracking and camera kept exposing. Simply re-opened gui and checked that dome was linked.
03/07/2022
Kurtis Williams
none
Andor
9.0
0
No problems to report
White dwarf time series photometry. I followed Brian's safe dome handling instructions - dome opened fully despite saying "ajar" and closed fully. Dome camera not responding, so used target images to verify closure. Great weather, 0.9" seeing. Used guider; it worked well (focus 975). Toward end of run dome stopped rotating at an azimuth of ~255 degrees. Cycled azimuth power, homed dome, re-linked it, and all was well.
03/06/2022
Kurtis Williams
none
Andor
0
10
Other
Did not open due to high stupidity (foggy brain, inability to focus on calendar)
03/03/2022
Keefe Kamp
Gokuldass
CCD
0
11
Humidity
TFOP and Be star observations were planned however to do a high chance of rain and humidity no observing took place
02/28/2022
Todd Hillwig
None
Andor CCD
0
10
Wind
High winds all night.
03/02/2022
Todd Hillwig
None
Andor CCD
0
10
Wind
High winds all night combined with last minute on-campus observation projects with students (we've had very few clear nights so far this semester).
02/24/2022
Tom Rutherford
None
CCD
0
9
Humidity
Unable to open due to high humidity.
02/23/2022
Ken Rumstay
None
CCD
0
11
Humidity
AGN's
02/22/2022
Todd Vaccaro
Nicolette Terracciano
Andor CCD
0
10
Equipment Failure
Eclipsing binary star. Controls for both the telescope and camera computers did not display when using Radmin. A reset of both computers was performed but still didn't display the controls. Did another reset for ACE (we thought maybe is was a connection issue, not a computer issue) and that displayed the controls. Once the controls were up, we couldn't open the dome. We tried following the dome door reset directions many times but the dome wouldn't open. Brian later logged on to help us. After his attempt, he informed us that the dome was stuck and could not be opened tonight for observing.
02/20/2022
Ted von Hippel
none
imaging
0
6
No problems to report
Decided not to make this night available to my students because of dome slit issues. Put out a call to SARA-Board@lists.fit.edu to see if anyone else wanted the night and no takers.
02/18/2022
Todd Vaccaro
Nicolette Terracciano
Andor CCD
3
8
Humidity
Eclipsing binary; missed primary eclipse due to humidity/clouds. Lost an hour early and ability to get flats b/c dome was not behaving; followed directions for door and Az issues, but had to restart services and cycle things as well. Ace (Tele) crashed on us 2 or 3 times, like once per hour.
02/17/2022
Ted von Hippel
Elvis Santana
imaging
0
6
Equipment Failure
The dome door opened but kept reporting it was in an opening status. We tried to close it and could not. I believe we followed the procedures to deal with the recent dome problems, but they failed for a number of attempts. After 3 or 4 hours, the same process (shut down power at the PDU, restart PDU power, hit the red octagon, then click on dome close) finally worked as seen on the camera within the dome.
02/16/2022
Xianming Han
none
CCD
10
0.5
Equipment Failure
eclipsing binaries The telescope mysteriously crashed. Restarting the ACE services on both computers fixed the problem. The dome also stopped rotating while observing.
02/15/2022
Xianming Han
none
CCD
10
0
Equipment Failure
eclipsing binaries the dome stopped rotating at the beginning of the night, and ACE helped . To fix dome not rotating problem, using the PDU to cycle the power to the Azimuth (be sure that its normal state is "OFF").
02/14/2022
Xianming Han
none
CCD
5
5
Humidity
eclipsing binaries humidity and the dome stopped rotating.
02/13/2022
Tom Rutherford
None
CCD
10
0
No problems to report
M31, open clusters, asteroid. No problems.
02/10/2022
Yong Sheng
Dieter Hartmann
RM griz
0
12
Equipment Failure
The dome can't be opened. Contacted Peter and he said that this issue hasn't been fixed yet
02/11/2022
William Keel
3 cats
Andor CCD
12
0
No problems to report
Dead quasar survey, H II regions in galaxies. Focus 96700, guider focus 900. Brian Brondel gave hints that made guiding much more reliable: use 2x2 binning and auto ROI instead of manually selecting a guide stars. I rewrote the guiding instructions and sent to the SARA mailing lists.
02/06/2022
Van Hamme
Samec
CCD
10
0
No problems to report
Observation of interacting binaries. AH Gem.
02/08/2022
Van Hamme
Samec
CCD
0
10
Other
Ice covering dome did not allow opening.
02/09/2022
Todd Vaccaro
Nicolette Terracciano and Aaron Brown-O'Dell
Andor
0
10
Clouds
Eclipsing binaries
02/06/2022
Roi Alonso
Lisa Altinier
CCD
9
2
Other
WD photometry, no autoguiding. Some problems to open the dome at the beginning: the "smart dome" switch at the Dome PDU had been set to "off". The focusser was not moving either, and required a full restart. After that, only the known issues (dome az reset was needed at some point). Last hour of the night the target had set.
02/05/2022
James webb
none
CCD
0
8
Other
senior care prevented me from observing. Offered it up but no one took it to my knowledge,
02/03/2022
Ken Watanabe
Nick Rimbert
N/A
0
10
Humidity
None.
01/30/2022
CCD
0
10
Equipment Failure
Dome would not open. Made numerous attempts, reset multiple times using the power outlet-- to no avail. Displayed "opening" and "closing," messages which never went away. Looking at it using the interior camera showed that the dome shutter was not responding to commands.
01/28/2022
Todd Vaccaro
Nicollette Terracciano
Andor
0
11
Humidity
eclipsing binaries. Never opened due to high humidity and threat of rain.
01/24/2022
William Keel
3 cats
Andor CCD
1.5
10
Humidity
Dead quasar survey. Had time to troubleshoot guiding a bit with Brian Brondel. BVR focus 95843, guider focus close to 0.
01/23/2022
Nuria Torres-Alba
Dieter Hartmann
SARA ORM
0
11
Humidity
No observations due to 100% humidity all night long
01/21/2022
Tom Rutherford
None
CCD
3
1
Software/Control PC Issues
Took several restarts before things started working properly-- with each restart, another function would come back online. Eventually, all was working and had no further issues.
01/18/2022
Tom Rutherford
None
CCD
0
4
Humidity
Unable to open due to high humidity the entire night
01/13/2022
Tom Rutherford
None
CCD
2
0
Other
M31. No real issues, except that the "dome closing" message still persists and never goes away. I did a couple of recycles of the shutter door power, but that did not seem to have any effect.
01/10/2022
Tom Rutherford
None
CCD
0
4
Humidity
Had to close just after getting flats due to high humidity and low temperatures. Dome never did give a "closed" message, but kept indicating "closing."
01/09/2022
William Keel
3 cats
Andor CCD
3
10
Clouds
Dead quasar survey. Focus 99150. Guider focus -600. Guider did new weird thing - offsets always showed 0 (more exactly listed number was something E-301), even when star disappeared behind a loud after guiding started..
01/07/2022
Yong Sheng
Dieter Hartmann
RM griz
0
13
Other
Observatory closed due to volcanic activity.
01/06/2022
Kurtis Williams
none
Andor
5.5
6.5
Wind
Deep H-alpha imaging lost due to high winds at start of night. White dwarf time series photometry in second half of night. No major problems. Winds were high but below limits for 2nd half of night. Seeing was marginal (2 arcsec or so); other ORM observatory DIMMs reported bad seeing too. Suggestion: Have wind speed in same units on Clarity, within the ACE interface (which I really liked tonight), and on the Closing limits on the SARA website. It's tough to convert m/s and kph to mph at 2am. Nice to be observing with this instrument again.
12/22/2021
Yong Sheng
None
RM
0
12
Other
Observatory closed due to volcanic activity.
12/14/2021
Kurtis Williams
none
Andor
0
13
Other
Surtr is preparing for Ragnarok and the battle against Aesir. Allegedly.
12/11/2021
Alex McDaniel
None
Andor CCD
0
12
Other
Observatory closed due to volcanic activity.
12/09/2021
Ken Rumstay
None
CCD camera
0
12
Other
Observatory closed due to volcanic activity.
12/05/2021
Martha Leake
none
Andor
0
10.8
Other
Asteroid light curves but volcanic eruptions continued with some high winds early and clouds. Leaving LP midnight.
11/16/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
0
12
Other
The observatory closed due to volcanic eruption
11/22/2021
Nuria Torres-Albà
Dieter
SARA ORM
0
12
Other
We could not observe due to the active volcano.
11/12/2021
Kurtis Williams
none
Andor
0
12
Other
Observatory closed due to volcanic eruption. Suggestion: GUI with option to turn plate tectonics off or on.
11/13/2021
Kurtis Williams
none
Andor
0
12
Other
Observatory closed due to volcanic eruption
10/09/2021
Kurtis Williams
nobody
Andor
0
12
Other
Was hoping to do H-alpha imaging, but was stymied by telluric silicate dust emission.
10/06/2021
Ted von Hippel
students
imaging
0
12
Other
Not opening due to volcanic ash risk.
09/03/2021
XIANMING HAN
none
CCD
6
0
No problems to report
eclipsing binary
09/03/2021
XIANMING HAN
none
CCD
4.5
0
No problems to report
eclipsing binary, sharing the night
09/02/2021
XIANMING HAN
none
CCD
8
0
No problems to report
eclipsing binary
09/01/2021
XIANMING HAN
none
CCD
7
0
No problems to report
eclipsing binary
09/22/2021
Tom Rutherford
None
CCD
0
8
Clouds
Overcast, high humidity.
09/20/2021
Roi Alonso
Lisa Altinier
Ccd
9.5
0
No problems to report
WD photometry. Avoided HA larger than 3h, and problems with tracking.
09/21/2021
Tom Rutherford
None
CCD
9
0.25
Equipment Failure
M31, Asteroid. Had to do a single restart at an hour angle of around +3 (not unexpected), but no issues other than that. A mostly uneventful night (which is a very good thing). No issues from the volcanic eruption.
09/18/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
8
0
Other
Asteroid photometry. No problems to report
09/18/2021
Núria Torres-Albà
Dieter Hartmann
SARA ORM
7
2
Software/Control PC Issues
Clemson University observations of high-z blazars At the beginning of the night, software crashed when trying to observe using a filter. Sequence tool message was: 'Filter Wheel Failed'. After that, software crashed. We did a full restart, and the same issue occurred. We contacted Peter and Brian, who fixed the issue remotely. The rest of night was smooth.
09/16/2021
Alex McDaniel
Dieter
Andor CCD
0
10
Humidity
Blazar Photo-z measurements. Unable to open/observe due to bad conditions (humidity/rain, wind, dust).
09/14/2021
Todd Hillwig
None
Andor CCD
9.5
0.5
Software/Control PC Issues
I could not control either the telescope or camera computers with either VNC or Radmin. Brian Brondel rebooted both computers remotely and they worked well after that.
09/12/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
0
8
Other
Asteroid photometry. The humidity changed over the night and reached the limit values. Beyond that, I had connections problems. I did connect on the computers, but these did not respond to my commands.
09/11/2021
Roi Alonso
Lisa Altinier
ccd
6
2
Equipment Failure
WD photometry. Tracking issues after the first half of the night, getting more and more frequent until we surrendered (tracking failing only few minutes after a full restart sequence).
09/11/2021
Saida Caballero
Fowler, Gokuldass, Van Putten
CCD
1.5
10
Equipment Failure
Massive star survey. Telescope would not accept slew command and returned a "Telescope unable to slew due to equipment failure." Moving the telescope with the hand paddle seemed to work but was not practical. Tried a simple restart and still had the same error.
09/09/2021
Judith Provencal
none
Andor
0
9
Other
too dusty to open
09/09/2021
Andy Baran
Tomomi Otani
Andor
9
1
Equipment Failure
1hr lost due to telescope failure, (frequently stops tracking after HA=2)
09/07/2021
Andy Baran
Tomomi Otani
Andor
9
1
Equipment Failure
The telescope keeps stopping after HA=3hrs again.
09/06/2021
Andy Baran
Tomomi Otani
Andor
9
1
Equipment Failure
The telescope keeps stopping after HA=2hrs, and this issue becomes very severe after HA=3hrs. Cycling ACE control cabinet power helps to get the telescope back in tracks but this task gets very irritating when one needs to do this every minute or so.
09/04/2021
Andy Baran
Tomomi Otani
Andor
8
0
Equipment Failure
The same tracking issue happened at HA=2hrs. The telescope just stopped tracking and cycling the ACE cabinet power fixed the issue. The camera was taking images even though the D drive was full and the images were not saved. No notification showed up, so be careful.
09/03/2021
Andy Baran
Tomomi Otani
CCD
6
2
Clouds
Cloud cover in the evening and during first 2 hours, the telescope stopped tracking right after HA=2hrs. restarting all ACE (including control cabinet) helped.
08/29/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
3
5
Equipment Failure
Asteroids photometry. Problems with the tracking after few hours. I had to fully restart the software a few times, but it did not fix the problem.
08/29/2021
Saida Caballero
Norton, Gokuldass, Kamp
CCD
6
6
Other
B star monitoring Dust counts were too high at the beginning of the night.
08/29/2021
Tom Rutherford
None
CCD
8
0
No problems to report
M31, Landolt fields. No problems at all-- everything worked well.
08/27/2021
William Keel
2 cats
Andor CCD
8
2
Software/Control PC Issues
Galaxies in H-alpha. Focus=101500, guider focus=1000. Initially could not control QSI guide camera. Brian and Peter got it connected, but then nonsense offsets were shown on the graph and any guide move stopped the drive (additional drive stops at all declinations around 3 hr west). Late in night, saw inconsistent count rates between frames even when GTC camera did not show clouds.
08/25/2021
Lisa Altinier
Roi Alonso
CCD
7
2
Equipment Failure
White Dwarf photometry. Tracking problems from the middle of the night.
08/25/2021
Alex McDaniel
Dieter Hartmann
Andor CCD
8
2.5
Equipment Failure
Blazar photo-z measurements. Technical issues: Had issues with tracking, and ultimately had to restart the software a few times to get things working. On one reset the node service wouldnt shutdown/startup properly but Paul and Brian helped resolve this. After that no other issues for the remainder of the night.
08/20/2021
Todd Vaccaro
None
Andor CCD
0
8
Other
Never opened due to dust. NOT weather page never fell below 30, but there were several domes open throughout the night including NOT. The site map never showed NOT had dust warning, only GTC showed it. The mirror cover was showing as open when I logged on and not knowing if there's an issue there I left as is. I only took bias frames.
08/18/2021
XIANMING HAN
none
CCD
2
6
Other
1. Dust. 2. Telescope keeps loosing tracking. DEC was 46 degrees, air mass ~1.2 when this happenned.
08/17/2021
Ken Rumstay
None
CCD
0
9
Other
Too much dust in the air (24-38 micrograms per cubic meter) to open dome. Gave up an hour before sunrise.
08/15/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
3.5
4.5
Equipment Failure
We observed about 3.5 hours until the tracking start to fail. I did the full restart two times, but it is not fixed the problem.
08/14/2021
Lisa Altinier
Roi Alonso
CCD
3
5
Equipment Failure
WD photometry. After about 3h of observations, tracking started to fail, and became unbearable shortly after, lasting only a few minutes after RA/Dec resets, and same after full restarts.
08/14/2021
Martha Leake
none
Andor
4.5
4
Clouds
Asteroid photometry. Dust at 27-23 at start of night, stable most of evening; after I closed b/c of clouds, aerosols increased to 35 suddenly. Clouds interfered for at least half of my 4 hours. No tracking problems, but incomplete slew ("unable to complete slew...") message forced a full restart, and a re-setting my protocols. File space disappeared without warning, so all asteroid files and later darks were lost. Slewing was fine after restart.
08/12/2021
Ken Watanabe
Nicolas RImbert
CCD
3
0.5
Equipment Failure
Blazars. Stopped tracking at the end.
08/11/2021
Todd Hillwig
None
Andor CCD
7
1
Equipment Failure
At the beginning of the night telescope tracked for about 30 minutes without issue on a target while using autofocus and waiting for twilight to end. Within 10 minutes of starting guiding, tracking turned off. After full restart didn't use guiding and tracking only turned off two more times (both very close to airmass 2 in the west, though).
08/10/2021
Ken Watanabe
Nicolas Rimbert
CCD
6
1
Equipment Failure
BL Lac and other Blazars. Telescope stopped tracking.
08/09/2021
Michael Fauerbach
NA
CCD
6
3
Equipment Failure
Asteroid photometry The first 3-4h were great, the telescope pointed and tracked fine, but then it stopped tracking. The usual reboot fixed that, but only for a short while.The period between the loss of tracking became shorter and shorter. After 5 reboots, it would not even track for a single exposure.
08/08/2021
XIANMING HAN
none
CCD
4
4
Equipment Failure
eclipsing binary, jupiter The telescope lost tracking often. It lost tracking, but displays it was still tracking, and could not slew. Only full restart will restore tracking.
08/07/2021
Yong Sheng
Dieter Hartmann
RM griz
7
3
Equipment Failure
The dome didn't move as the telescope moves. The tracking was also done.
08/06/2021
Stephen Gillam
None
Andor camera
5
5
Software/Control PC Issues
The system worked well until we got to about 1hr and 20 mins west of the meridian. The scope was at an airmass of 1.12 and an altitude of ~62 deg. It was no where near any of the limits. Then, the scope suffered the first of five episodes of tracking failure and/or the "unable to slew due to equipment failure". In the first two instances, reselecting the target and slewing (the short distance) to it was sufficient to get it tracking again. Over the next three hours the tracking failed three more times with the "..due equipment failure" message. I notices that the not only was the RA tracking not working, but the the scope was also drifting in declination. These episodes were fixed by doing the full ACE restart procedure. Except that the filter wheel GUI on the camera computer also stopped working during the fourth episode. This resulted in a loss of filter and other info. in the file headers. During the fifth instance, the filter wheel GUI went away and could not be coaxed into starting up again. I tried two full restarts for this one. I was able to get bring the telescope control up again and park the scope. During two of episodes the camera GUI froze. It displayed the following error message. "tr1::bad_weak_ptr." Even after the full ACE restart procedure the old instance of the ACE control system was still present. It would crash and go away only when I clicked buttons on the service selector. The, a new instance of the camera control could be started.
08/05/2021
Todd Hillwig
None.
Andor CCD
2
7
Equipment Failure
Nearly complete waste of a beautiful night. Tracking turned off repeatedly. It took four complete restarts to finally get the telescope to keep tracking for more than 10 or 15 minutes. Observed for 2 hours and tracking turned off again, repeatedly. Three more complete restarts and I finally gave up. Tried several positions on the sky, with no success. Even when it was tracking it didn't track well. With the guiding on (which it didn't seem was very effective), the telescope drifted by several arc seconds every ten minutes or so. Also, the "Park" entry under "Presets" sends the telescope to the wrong position: it sends it to an RA of 0:00:00 NOT an HA of 0:00:00...and leaves the telescope tracking, which is obviously very bad. Oh, and the mirror covers never read closed, still read "closing."
08/03/2021
Roi Alonso
Lisa Altinier
CCD
3
5
Wind
First half of night with high wind gusts. Second part ok, with already known problems. Tracking was difficult/impossible at end of night, too many RA resets needed. Not clear status of mirror petals, they seem to be open, but closing them takes forever and seems not to have an effect (opening again is almost immediate).
08/04/2021
Tom Rutherford
None
Andor CCD
8
0.5
No problems to report
M31, Kepler exoplanet. Had to do three restarts at around 2:30 UT but that is fairly typical behavior so no real issues to report-- an uneventful night.
07/30/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
6
2
Other
I began the observation night very well without problems. In the middle of the night, it was necessary to do the full restart because the slew shows me an error (unable to complete the slew due to an equipment error), I did a full restart again and it was fixed. The big problem was at the end (close to 03:00 UT), it was necessary to do the full restart two times, and it was impossible to continue the observation. I finished because the tracking did not work.
08/02/2021
Todd Vaccaro
N/A
Andor CCD
5
3
Equipment Failure
Eclipsing binaries. Missed early flats due to telescope RA not reporting correctly to fits headers, which apparently is only occurring when tele is idle. Mirror cover was saying "closing" when I first logged on and continued to do so after I closed. I lost the majority of an eclipse egress due to tracking failures (over 5 times before I gave up on the target). It was failing ~+2 hrs HA at +78 deg. A full reset was needed to move the tele anywhere as it kept reporting an equipment error. After a full reset I could reacquire the target, but it would produce trailed stars after a few frames. When I went to another object that had yet to transit it started doing the same thing at HA 00:15:00. Something wrong on W side of mount?
07/30/2021
Keefe Kamp
None
CCD
0
10
Other
Cyg X-1 and Be Star Observations Planned but was too dusty to open
07/29/2021
Keefe Kamp
None
CCD
5
1
Equipment Failure
Cyg X-1 and Be star observations. Problems with petals. where it would continuously say closing but would not close. Also when telescope starting tracking it would change from open to opening but everything appear normal. Full restart fixed both issues
07/28/2021
Núria Torres-Albà
Dieter Hartmann
SARA ORM
6.5
1
Software/Control PC Issues
Clemson University observations of high-z blazars. Problems: -Computer PC was unresponsive at the beginning of the night. Called Brian and he rebooted it. -Tracking stopped working (could see star traces on the image). Stopped tracking and attempted to restart it, but 'Slew Telescope' window showed an error (Unable to complete slew due to an equipment error). Restarted the system as usual (Full Restart). Problem persisted. -After restart 'Mirror Cover' window was stuck in 'Opening' for almost 5min. -Restarted once more. This time, no issue with mirror covers. Problem was only fixed for one single exposure (5min) and then returned. -Restarted a third time, and switched to a higher altitude source. Tracking worked afterwards.
07/27/2021
Yong Sheng
None
RM griz
0
10
Wind
Clemson University blazar study. The wind speed is over or around the threshold.
07/25/2021
Todd Hillwig
None
Andor CCD
9
0
Other
At the start of the night when I opened the dome, it read as opening, but never got to "Opened." Restarting the software fixed this. The telescope tracking was not great. Five minute exposures showed drifting. There is no option to open the Guider GUI, so I couldn't guide. When I sent the telescope to Park position, it went to the correct place then started tracking. Had to hit Cancel to get it to stop tracking. And after closing the dome, it says "Doors ajar."
07/22/2021
Roi Alonso
none
CCD
4
4
Clouds
WD photometry, no filter, focus around 103400. Full restart needed to be able to move the telescope at the beginning of night. Telescope stopped tracking around middle of night, but conditions were cloudy anyway.
07/18/2021
Robert Berrington
Todd Vaccaro
Arc Camera
0
9
Equipment Failure
Observatory closed due to mechanical failure...Did not open.
07/21/2021
Tom Rutherford
None
Andor CCD
6
2
Equipment Failure
H-alpha filters were missing (this was the most critical issue I had). Dome did not point with telescope at first, but got them re-synced. CCD computer gave only a black screen with no controls visible-- eventually came back, but was unresponsive (as if it were only an image of the screen instead of the screen itself)-- I was able to control the CCD through the Telescope screen. Usual multiple restarts when changing targets later in the night.
07/16/2021
Javier Licandro
Hissa Medeiros
Andor camera (JKT)
0
8
Equipment Failure
Equipment Failure. The problems with the mirror cover are not fixed.
07/16/2021
Martha Leake
none
Andor
1
7
Equipment Failure
Asteroid photometry. Was able to get biases, some darks, and flats (to be transferred later), but unable to get any stellar images because of problems with the mirror cover. I didn't fully know that it was fixed, but apparently was able to get light to the camera for my flats. The mirror cover never fully opened, but did shut down quickly. Dome azimuth was checked and improved with the Dome Azimuth restart. Ken Rumstay, Brian Murphy and Peter Mack all helped on various pieces of the problem. I was/am embarrassed that this was akin to forgetting to remove the camera lens cover. But the problem is tougher and will be looked at Monday. Good luck to all involved. (Dust was not a problem.)
07/12/2021
Roi Alonso
none
ccd
0
8
Clouds
As in previous reports, sky was cloudy and dust above 40 ugr/m3 as reported in the tng weather server. Did not try to open.
07/12/2021
James Webb
none
CCD
0
10
Other
Blazar microvariability. Lots of dust and sky was obscured on all sky cameras at the start of the night so I threw in the towel.
07/11/2021
Keefe Kamp
None
CCD
0
10
Clouds
Cyg X-1 and Be stars planned but was too dusty and cloudy to observe
07/10/2021
Ken Rumstay
None
CCD camera
5
3
Other
AGN monitoring program. No mechanical or software problems to report. Telescope pointing and tracking was excellent. High aerosol levels at start of night and then again just before morning twilight.
07/10/2021
Stephen Gillam
Thiyagarajan, Alamos, Nelson
Andor camera
0
10
Equipment Failure
r0Ap Telescope unable to slew due to equipment failure. This is presumed to be due to driving the scope to the west limit, reported last night (7/8 -7/9/21).
07/09/2021
Gillam
Pepe, Alamos, Thiyagarajan, Nelson
Ansor
10
0
Other
roAp stars. At the start of the night, the camera and telescope computers were locked up, We power-cyc;led them. The telescope control was then not talking to the hardware. We carried out the procedure to restart the ACE node services and the camera and telescope controls then worked. Shortly after this, the dome stopped moving. We recycled the Azimuth control and homed the dome. The dome the worked perfectly until about an hour before the end of the night. The dome again needed to azimuth-cycled and homed. This was successful. I then allowed the scope to reach the western limit. It is currently stuck there. Could you move the scope off the limit? - Thanks!
07/07/2021
Stephen Gillam
Lilyanne Pepe, Leila Alamos, Chloe Nelson, Aakash Thiyagarajan
Andor Camera
10
0
No problems to report
Image type header keyword is missing from FITS headers. roAp stars
07/06/2021
Xianming Han
none
none
0
8
Software/Control PC Issues
ACE was upgrading Windows 10, and after the upgrade, something went bad with the telescope.